The subjects in the data are college students. In the data, id is student ID, anxiety is student’s anxiety score via Anxiety Scale, selfest is student’s self-esteem score via Rosenberg Self-esteem Scale, GPA is student’s GPA; for gender, 0=female, 1=male; for grade, 1=freshman, 2=junior, 3=senior. We have known that population mean for Anxiety Scale is μ=60 with σ=10.

Raise relevant questions ( 2 questions is fine) about the data extensively, the questions can be either about descriptive analysis or inferential analysis, for example, “what are the mean and standard deviation of …” or “whether female students are more … than male students”. Then you should select appropriate statistical analyses for your questions and explain why you select these analyses. Finally, conduct analyses and report the results

This is the data

id anxiety selfest gender grade GPA
1 61 25 0 1 3.31
2 71 29 0 1 2.67
3 67 28 1 1 3.89
4 52 28 0 1 4.00
5 81 20 1 1 2.00
6 81 21 1 1 2.56
7 69 28 1 1 3.12
8 60 38 0 1 3.47
9 67 29 1 1 3.29
10 58 28 0 1 3.22
11 48 28 1 1 3.68
12 66 24 1 1 3.78
13 49 37 1 1 3.84
14 75 25 1 1 3.11
15 79 27 1 1 3.05
16 84 21 0 1 3.33
17 70 22 1 1 2.21
18 52 35 0 1 2.95
19 63 33 0 1 2.44
20 62 32 1 1 3.00
21 70 30 1 1 3.91
22 63 25 1 1 2.44
23 61 37 0 1 3.68
24 71 28 0 1 3.27
25 68 29 1 1 3.18
26 62 36 0 1 2.47
27 64 28 0 1 3.45
28 62 28 1 1 3.67
29 63 33 1 1 3.78
30 60 25 0 1 3.79
31 57 29 0 1 3.10
32 57 35 0 1 3.20
33 80 23 0 1 2.10
34 68 28 0 1 2.89
35 70 26 1 1 2.71
36 58 28 0 1 3.02
37 73 24 0 1 3.09
38 65 35 0 1 3.94
39 51 36 0 1 3.95
40 72 23 1 1 3.03
41 62 29 1 1 3.06
42 76 31 1 1 2.94
43 68 30 1 1 2.86
44 70 32 1 1 2.06
45 65 30 0 1 3.14
46 70 28 1 1 3.19
47 35 34 0 1 3.29
48 54 30 0 1 3.64
49 62 30 0 1 3.49
50 66 31 0 1 2.78
51 79 24 0 2 1.67
52 86 18 0 2 3.18
53 65 33 1 2 4.00
54 64 35 1 2 3.97
55 78 32 0 2 3.25
56 71 25 1 2 3.54
57 65 27 1 2 3.75
58 49 41 0 2 2.22
59 70 26 1 2 2.68
60 74 27 0 2 3.45
61 60 35 1 2 3.54
62 60 34 0 2 3.65
63 71 24 1 2 3.62
64 62 38 0 2 3.69
65 60 32 1 2 2.99
66 85 26 0 2 3.07
67 69 29 0 2 3.91
68 51 35 1 2 3.89
69 62 35 1 2 3.23
70 69 19 1 2 3.36
71 44 38 1 2 3.61
72 67 27 1 2 3.04
73 78 32 0 2 2.91
74 53 29 0 2 2.59
75 80 22 1 2 3.03
76 62 34 1 2 3.56
77 74 27 0 2 3.67
78 64 25 0 2 3.85
79 76 20 1 2 3.88
80 44 33 1 2 3.47
81 60 27 1 2 3.22
82 69 25 1 2 3.03
83 74 25 1 2 3.09
84 95 19 1 2 2.89
85 75 27 0 2 2.67
86 77 20 1 2 3.01
87 76 21 1 2 3.01
88 58 31 0 2 3.22
89 53 33 1 2 3.24
90 56 27 0 2 3.53
91 73 31 0 2 3.64
92 64 38 0 2 3.71
93 69 30 0 2 3.89
94 57 33 1 2 3.01
95 92 23 1 2 2.10
96 75 24 0 2 2.19
97 97 24 0 2 2.52
98 74 28 0 2 3.09
99 59 31 0 2 3.03
100 62 23 0 2 3.45
101 47 38 0 3 3.99
102 60 26 0 3 4.00
103 68 23 0 3 4.00
104 67 32 1 3 3.89
105 50 31 0 3 3.78
106 76 24 1 3 3.67
107 56 35 1 3 3.56
108 59 29 1 3 3.67
109 72 27 0 3 3.88
110 42 37 1 3 3.81
111 45 33 0 3 3.17
112 67 30 1 3 3.67
113 59 34 0 3 2.22
114 55 37 1 3 3.98
115 52 28 0 3 2.89
116 84 21 1 3 3.78
117 65 27 0 3 2.87
118 50 33 1 3 3.67
119 82 27 0 3 2.99
120 78 22 1 3 3.78
121 59 27 0 3 3.01
122 75 31 1 3 3.86
123 91 23 0 3 3.57
124 58 27 0 3 3.66
125 67 32 0 3 3.89
126 59 28 0 3 2.62
127 61 31 0 3 3.05
128 44 32 1 3 3.01
129 67 16 0 3 3.44
130 62 28 0 3 3.47
131 66 31 1 3 3.89
132 68 30 0 3 3.87
133 63 34 1 3 3.78
134 79 22 1 3 2.55
135 60 29 0 3 3.03
136 67 31 1 3 3.09
137 71 32 1 3 3.95
138 59 36 1 3 3.78
139 78 24 1 3 3.59
140 56 32 1 3 3.66
141 66 31 1 3 3.70
142 68 30 1 3 3.78
143 72 37 0 3 2.99
144 75 29 0 3 3.98
145 74 27 0 3 3.27
146 78 22 1 3 3.49
147 63 31 1 3 3.50
148 90 14 1 3 3.61
149 42 32 1 3 3.88
150 62 32 1 3 3.93
Answer #1

Solution

I Descriptive Analysis of GPA

Back-up Theory

Let xi = ith value, i = 1, 2, ……., n. Then,

Mean (Average), µ, = (1/n)Σ(i = 1, n)(xi) ………………………………………………………………..……………………. (1)

Variance, σ2 = (1/n)Σ(i = 1, n){(xi – µ)2} or equivalently [{(1/n)Σ(i = 1, 20)(xi)2} – µ2 ] …………………... (2)

Standard deviation (SD), σ = sqrt(Variance) …………………………………………………………………………..….. (3)

Median

Let x(1) , x(2) , x(3) , ……….. , x(n - 1) , x(n) be the ordered set of the given values; i.e.,

x(1) < x(2) < x(3) < ……….. < x(n - 1) < x(n)

Case 1: n is even, say n = 2k

Median = Average of two middle values in the ordered set ……………………………………………………… (4a)

= (x(k) + x(k + 1))/2

Case 2: n is odd, say n = 2k + 1

Median = (k + 1)th value in the ordered set; i.e., x(k + 1) ……………………………………………………… (4b)

First Quartile Q1 is the median of the first half of the ordered set ……………………………………….. (5a)

and Third Quartile Q3 is the median of the second half of the ordered set…………………………….. (5b)

Now, to work out the solution

Vide (1), Mean = 3.3067 Answer 1 [using Excel Function Statistical AVERAGE]

Vide (3), Standard deviation = 0.5146 Answer 2 [using Excel Function Statistical STDEVP]

Vide (4a), Median = 3.345 Answer 3 [Average of 75th and 76th observations in the ordered set – given at end]

Vide (5a), Q1 = 3.01 Answer 4 [38th observations in the ordered set – given at end]

Vide (5b), Q3 = 3.75 Answer 5 [113th observations in the ordered set – given at end]

Minimum value = 1.67 Answer 6 [1st observations in the ordered set – given at end]

Maximum value = 4.00 Answer 7 [150th observations in the ordered set – given at end]

II Inferential Analysis of Selfest

Two sample t-test for equality of means between male and female.

Let X = selfest score of male

      Y = selfest score of female

Then, X ~ N(µ1, σ12) and Y ~ N(µ2, σ22), where σ12 = σ22 = σ2, say and σ2 is unknown.

Claim:

Male and female do not differ with respect to selfest score

Hypotheses:

Null: H0: µ1 = µ2 Vs Alternative: HA: µ1 µ2

Test Statistic:

t = (Xbar - Ybar)/[s√{(1/n1) + (1/n2)}] where

s2 = {(n1 – 1)s12 + (n2 – 1)s22}/(n1 + n2 – 2);

Xbar and Ybar are sample averages and s1,s2 are sample standard deviations based on n1 observations on X and n2 observations on Y respectively.

Calculations

Summary of Excel calculations is given below:

n1 =

77

n2 =

73

Xbar =

28.41558

Ybar =

29.28767

s1 =

5.289888

s2 =

4.820326

s =

5.066891

tcal =

-1.05361

α =

0.05

tcrit =

1.976122

p-value =

0.293779

Distribution, Significance Level, α Critical Value and p-value:

Under H0, t ~ tn1 + n2 - 2. Hence, for level of significance α%, Critical Value = upper (α/2)% point of tn1 + n2 - 2 and p-value = P(tn1 + n2 - 2 > | tcal |) = 2P(tn1 + n2 - 2 > tcal) if tcal > 0 and 2P(tn1 + n2 - 2 < tcal) if tcal < 0

Using Excel Function: Statistical TINV and TDIST, these are found to be as shown in the above table.

Decision:

Since | tcal | < tcrit, or equivalently since p-value > α, H0 is accepted.

Conclusion:

There is sufficient evidence to suggest that the claim is valid. ANSWER

DONE

GPA ordered set

id

GPA [min to max]

1

1.67

2

2

3

2.06

4

2.1

5

2.1

6

2.19

7

2.21

8

2.22

9

2.22

10

2.44

11

2.44

12

2.47

13

2.52

14

2.55

15

2.56

16

2.59

17

2.62

18

2.67

19

2.67

20

2.68

21

2.71

22

2.78

23

2.86

24

2.87

25

2.89

26

2.89

27

2.89

28

2.91

29

2.94

30

2.95

31

2.99

32

2.99

33

2.99

34

3

35

3.01

36

3.01

37

3.01

38

3.01

39

3.01

40

3.02

41

3.03

42

3.03

43

3.03

44

3.03

45

3.03

46

3.04

47

3.05

48

3.05

49

3.06

50

3.07

51

3.09

52

3.09

53

3.09

54

3.09

55

3.1

56

3.11

57

3.12

58

3.14

59

3.17

60

3.18

61

3.18

62

3.19

63

3.2

64

3.22

65

3.22

66

3.22

67

3.23

68

3.24

69

3.25

70

3.27

71

3.27

72

3.29

73

3.29

74

3.31

75

3.33

76

3.36

77

3.44

78

3.45

79

3.45

80

3.45

81

3.47

82

3.47

83

3.47

84

3.49

85

3.49

86

3.5

87

3.53

88

3.54

89

3.54

90

3.56

91

3.56

92

3.57

93

3.59

94

3.61

95

3.61

96

3.62

97

3.64

98

3.64

99

3.65

100

3.66

101

3.66

102

3.67

103

3.67

104

3.67

105

3.67

106

3.67

107

3.67

108

3.68

109

3.68

110

3.69

111

3.7

112

3.71

113

3.75

114

3.78

115

3.78

116

3.78

117

3.78

118

3.78

119

3.78

120

3.78

121

3.78

122

3.79

123

3.8

124

3.84

125

3.85

126

3.86

127

3.87

128

3.88

129

3.88

130

3.88

131

3.89

132

3.89

133

3.89

134

3.89

135

3.89

136

3.89

137

3.91

138

3.91

139

3.93

140

3.94

141

3.95

142

3.95

143

3.97

144

3.98

145

3.98

146

3.99

147

4

148

4

149

4

150

4